# Parameter-Efficient and Personalized Federated Training of Generative Models at the Edge

### v1

@@ 5.3 Effectiveness of Personalization
While a strong global model is desirable, the ultimate goal of pFL is to deliver superior performance for individual users. We evaluate this by testing the models on each client’s local, held-out test data. The results, averaged across all 100 clients, are summarized in Table LABEL:tab:personalization.
@@
The same trend is observed for CIFAR-10 image generation. The global FedGen-Edge model achieves an average local FID of 40.1. After one epoch of personalization, the average FID improves to 32.8. This personalized model significantly outperforms the ‘Local-Only‘ model (89.6 FID) and the generic global model. The visualization in Fig. LABEL:fig:personalization_dist, which shows a histogram of the performance gains (reduction in FID) from personalization, reveals that this benefit is widespread, with over 90% of clients experiencing a notable improvement. This confirms that the benefits are not limited to a few clients but are a systemic property of our framework. Compared to Ditto li2021ditto , which uses regularization to balance global and local objectives, our approach offers a more explicit and intuitive mechanism for personalization: learn a good global initialization collaboratively, then specialize it locally.
@@ 5.4.1 Impact of LoRA Rank
The rank ‘r‘ controls the capacity of the adaptation module. A higher rank allows the adapter to represent more complex transformations but also increases the number of trainable parameters and thus the communication cost. Fig. 6 plots the final global model performance (PPL and FID) as a function of ‘r‘. We observe a clear trend: as the rank increases from 2 to 32, performance generally improves. However, the improvements exhibit diminishing returns. For the PTB task, the most significant gains occur when moving from $r=2$ to $r=8$, with only marginal improvements thereafter. A similar pattern holds for CIFAR-10. This suggests that the "intrinsic rank" of adaptation for these tasks is indeed low, and a relatively small rank (like our choice of $r=8$ for PTB and $r=4$ for CIFAR-10) is sufficient to capture most of the necessary information. This is a crucial finding, as it allows us to choose a rank that balances performance and efficiency, a principle that could be extended to federated neural architecture search he2020fednas , 10251628 to find optimal, resource-aware configurations.
@@ 5.5 Qualitative Analysis and Visualizations
Quantitative metrics provide a measure of overall performance, but qualitative examples offer a more intuitive understanding of the models’ capabilities. In Fig. LABEL:fig:qualitative_images, we present a grid of images generated by different models for the CIFAR-10 task. The images from the Centralized model are sharp, diverse, and clearly recognizable. The images from our personalized FedGen-Edge model are of comparable quality. In contrast, the images from the global FedAvg model show clear artifacts and mode collapse, often failing to generate certain classes. The ‘Local-Only‘ models can only generate images from the classes they were trained on, resulting in a complete lack of diversity from a global perspective. These visualizations vividly illustrate the success of our method in preserving generation quality and diversity in a decentralized setting.
@@ 5.6 Discussion and Broader Implications
While our results are promising, we acknowledge the limitations of this work. We operated under an honest-but-curious server model and did not consider malicious clients. Future work should investigate the security of federated PEFT, as the low-dimensional updates might present a different attack surface for privacy inference zhu2019deep or backdooring bagdasaryan2020how compared to full model updates. Furthermore, while we demonstrated our framework on moderately sized models, scaling to models with hundreds of billions of parameters will require further system-level optimizations, potentially integrating techniques from decentralized learning liu2025accelerating and alternative paradigms like Split Learning vepakomma2018split . Finally, extending our framework to more complex, multi-modal generative tasks, such as those combining vision and language for emotion recognition 10149418 or video grounding hu2025unified , represents an exciting avenue for future research. The principles of FedGen-Edge, however, provide a strong and scalable foundation for this next generation of distributed, generative artificial intelligence.

### v3

@@ 5.3 Effectiveness of Personalization
While a strong global model is desirable, the ultimate goal of pFL is to deliver superior performance for individual users. We evaluate this by testing the models on each client’s local, held-out test data. The results, averaged across all 100 clients, are summarized in Table.
@@ 5.4.1 Impact of LoRA Rank
The rank ‘r‘ controls the capacity of the adaptation module. A higher rank allows the adapter to represent more complex transformations but also increases the number of trainable parameters and thus the communication cost. Fig. 6 plots the final global model performance (PPL and FID) as a function of ‘r‘. We observe a clear trend: as the rank increases from 2 to 32, performance generally improves. However, the improvements exhibit diminishing returns. For the PTB task, the most significant gains occur when moving from $r=2$ to $r=8$, with only marginal improvements thereafter. A similar pattern holds for CIFAR-10. This suggests that the "intrinsic rank" of adaptation for these tasks is indeed low, and a relatively small rank (like our choice of $r=8$ for PTB and $r=4$ for CIFAR-10) is sufficient to capture most of the necessary information. This is a crucial finding, as it allows us to choose a rank that balances performance and efficiency, a principle that could be extended to federated neural architecture search he2020fednas to find optimal, resource-aware configurations.
@@ 5.5 Qualitative Analysis and Visualizations
Quantitative metrics provide a measure of overall performance, but qualitative examples offer a more intuitive understanding of the models’ capabilities.we present a grid of images generated by different models for the CIFAR-10 task. The images from the Centralized model are sharp, diverse, and clearly recognizable. The images from our personalized FedGen-Edge model are of comparable quality. In contrast, the images from the global FedAvg model show clear artifacts and mode collapse, often failing to generate certain classes. The ‘Local-Only‘ models can only generate images from the classes they were trained on, resulting in a complete lack of diversity from a global perspective. These visualizations vividly illustrate the success of our method in preserving generation quality and diversity in a decentralized setting.
@@ 5.6 Discussion and Broader Implications
While our results are promising, we acknowledge the limitations of this work. We operated under an honest-but-curious server model and did not consider malicious clients. Future work should investigate the security of federated PEFT, as the low-dimensional updates might present a different attack surface for privacy inference or backdooring bagdasaryan2020how compared to full model updates. Furthermore, while we demonstrated our framework on moderately sized models, scaling to models with hundreds of billions of parameters will require further system-level optimizations, potentially integrating techniques from decentralized learning and alternative paradigms like Split Learning. Finally, extending our framework to more complex, multi-modal generative tasks, such as those combining vision and language for emotion recognition 10149418 or video grounding, represents an exciting avenue for future research. The principles of FedGen-Edge, however, provide a strong and scalable foundation for this next generation of distributed, generative artificial intelligence.